# Understanding The Concerns and Choices of The Public When Using Large Language Models for Healthcare

### v1

@@ 3.2. Interviews
The interviews were semi-structured. We started by asking about the basic background of the participants. We divided the interview questions into three parts: (1) motivations for using LLMs for healthcare information; (2) misinformation and privacy issues encountered when using LLMs; and (3) the role of LLMs in the doctor-patient relationship. A detailed list of questions is provided in Section 7.1 in the appendix.
After each interview, we held a brief meeting to discuss emerging themes. Two authors independently conducted open and axial coding of the interview transcripts (strauss1990qualitative), and regularly met to compare their results and discuss the coding process. If the two coders could not reach a consensus, a third researcher would join the discussion to help make a final decision. Emerging themes included …, …, …, …, and so on. In the paper, we use anonymized quotes to illustrate the main themes.
@@ 4.1. How The Public Obtain Healthcare Information? (RQ1)
LLMs are gaining popularity as a means of healthcare information seeking. Among the 167 valid surveys, 162 participants had used the Internet (including but not limited to search engines, online communities, AI-based healthcare chatbots, and LLMs) to search for healthcare information. Among them, 157 (97% ) had used search engines to search for healthcare information; 127 participants had obtained healthcare information from online communities, and 123 participants had obtained healthcare information from LLMs. LLMs like ChatGPT have only been around for about a year, but they have already become a frequently used way for the public to obtain health information. In contrast, only 62 participants had used Al-led healthcare chatbots which have existed for a longer time than LLMs. The percentage of users who used chatbots for healthcare Q&A was slightly larger than in a previous study (Nadarzynski et al., 2019), but chatbots were still not a mainstream way for the public to obtain healthcare information. We confirmed this during the interviews.
@@
There was more than one way to obtain healthcare information from the Internet, and they were often used together for better information outcomes. Of the 123 participants who had used LLMs to obtain health information, 105 participants cross-validated the healthcare information obtained from LLMs with healthcare information obtained from other sources. During the interviews, we found that due to the popularity of online consultation platforms in China (Ding et al., 2020; Hao, 2015; Hao et al., 2017), participants originally from China had heavily drawn on them for healthcare information. The students majoring in medicine (P3, P16) would also use more professional medical databases, such as the Merck Manual of Diagnosis and Therap. Other findings were largely consistent with the survey – the participants interviewed had all used search engines to search for healthcare information before, and most had obtained health information through online communities, but Among the 17 participants who had used LLMs to search for health information, only P5 and P9 had used AI-led chatbots.
@@
In the survey, based on some previous studies (Rice, 2006; Nadarzynski et al., 2019; Xiao et al., 2014), we roughly divided the health information obtained into four categories: (1) daily life health information (such as reasonable diet); (2) medication information (such as drugs and hospitalization); (3) symptom self-diagnosis (such as searching for disease); and (4) searching for health information again after being diagnosed by a doctor due to doubts about the treatment plan. Among the 160 participants who had searched for health information, 144, 145, 145, and 139 participants had searched for the above four types of information respectively. The search frequency for the four types of information was all very high and there was no significant difference. This phenomenon showed that the public was not only paying attention to diseases but also their healthcare in daily life. Similarly, during the interviews, we found that the participants mostly used ChatGPT to search for symptom-related information, but also used it to search daily health information such as reasonable diets and fitness plans. For example, P2 only uses Chatgpt to initially develop a fitness plan:
@@
In terms of frequency of searching for health information, 95 participants in the survey searched for health information once a month or more often; during the interview, most participants said they would search for health information every one to two weeks. A few interview participants like P1 searched for healthcare information nearly every day.
@@ 4.2. Motivation for Using LLMs for Healthcare Information (RQ2)
During the interviews, we asked the participants in detail about their motivations for using LLMs to obtain health information. We were particularly concerned about why participants chose to use LLMs like ChatGPT to obtain health information rather than other methods such as search engines, online communities, or consulting doctors online and offline.
Compared with search engines, using ChatGPT to obtain healthcare information was deemed faster and more accurate. Most of the participants were students. Before using ChatGPT to search for health information, they had used it to search for solutions to study problems. They believed that ChatGPT was more helpful to them than search engines in solving academic problems. This prompted them to start using ChatGPT for healthcare information as well.
@@
P5 also had the same idea. She believed that some subjective feelings of drug use cannot be described in detail by Chatgpt. ChatGPT rarely provided emotional support. Thus most participants expressed that they would only use it as a supplement or substitute for search engines, instead of as a chatting buddy.
Compared to offline or online medical consultations with doctors, the participants thought obtaining healthcare information from LLMs was more convenient and quick. Making an appointment with a doctor offline not only cost money but also took a lot of time; even if one made an appointment with a doctor online, there was no guarantee that the doctor would always be online to answer questions. However, all the participants acknowledged that if they felt their condition was serious and needed medical equipment for examination, they would still choose to go to the hospital. In other words, all participants would only use LLMs when they were experiencing minor discomfort or searching for daily healthcare information.
@@ 4.3.1. Misinformation and Inaccuracy of LLM Responses
LLMs were deemed more accurate than other information-seeking channels for healthcare information. In the survey, we designed two Likert-scale questions to probe this issue: “You can find more accurate health information with LLMs (like ChatGPT) than search engines or online communities” and “Compared with using search engines or online communities, using LLMs can reduce misinformation.” Excluding participants who were unable to answer these questions, e.g., those who had never used the Internet to search for health information, a total of 160 participants answered these two questions. For the first statement, 83 participants “agree” or “strongly agree” with it, accounting for 52% of the participants; 34 participants “disagree” or “strongly disagree” with it, accounting for 21% of the participants; The remaining participants neither agree nor disagree with it. For the second statement, 81 participants “agree” or “strongly agree” with it, accounting for 51% of the participants; 42 participants “disagree” or “strongly disagree” with it, accounting for 26% of the participants; the remaining participants neither agree nor disagree with it. As we can see, more than half of the participants agreed that LLMs are more accurate and produce less misinformation compared to other methods like search engines.
@@
During the interviews, we also asked participants how they judged misinformation, or how they prevented “hallucinations” from LLMs. Most of participants did not completely give up using search engines or online communities, and they would cross-verify the health information obtained from LLMs with health information obtained through other channels. Some participants, such as P12 and P14, would consult doctors after asking LLMs for health information, and their trust in LLMs was enhanced by the doctor’s feedback. However, there are exceptions. P8 said that since using Chatgpt, he has never used a search engine again because he hates the misinformation on Google.
@@ 4.3.2. Doctor-Patient Relationship
In the survey, we designed three questions to explore this question: “Patients using LLMs to self-diagnose before asking the doctor can improve the doctor’s diagnostic performance,” “I think doctors using LLMs to diagnose and treat conditions is reasonable” and “I think doctors using LLMs for medical assistance work (such as writing medical records) is reasonable.” A total of 160 participants answered these three questions. For the first question, 79 participants “agree” or “strongly agree” with the statement, accounting for 50% of the participants; 40 participants “disagree” or “strongly disagree” with the statement, accounting for 25% of the participants; the remaining participants neither agree nor disagree. For the second question, 75 participants “agree” or “strongly agree” with the statement, accounting for 47% of participants; 43 participants “disagree” or “strongly disagree” with the first question above, accounting for 27% of the participants. For the last question, 104 participants “agree” or “strongly agree” with the statement, accounting for 65% of the participants; 25 participants “disagree” or “strongly disagree” with the statement, accounting for 16% of the participants; the remaining participants neither agree nor disagree. It can be seen that most people thought it reasonable for doctors to use LLMs to perform auxiliary work such as writing medical records. When it came to using LLMs for diagnosis or self-diagnosis, the participants were more doubtful.
During the interviews, most of our participants believed that self-diagnosis assisted by LLMs could help doctors better judge their conditions. This is inconsistent with the survey. P6 talked about the importance of knowing one’s own medical history and physical conditions before seeing a doctor:
@@
A few participants believed that LLMs could be used for diagnosis since AI was an unstoppable trend and would eventually replace doctors. P7 noted:
@@ 5. Discussion
In this section, we discuss recommendations for the public to use LLMs to obtain health information and conduct self-diagnosis and recommendations for doctors to use LLMs to diagnose patients. We further give recommendations for future research in medical LLMs.
@@ 5.1. Recommendations for The Public
The medical field is rapidly developing – academic results that are correct today may be overturned or proven to have certain limitations tomorrow (Singhal et al., 2023). Moreover, since everyone’s physical condition is different, the same symptoms may come from different diseases for different people. We ought to fully consider these medical backgrounds and make contextual recommendations for the public regarding using LLMs to obtain health information to maximize public interests.
According to our results, although LLMs cannot completely solve the problem of misinformation generated by search engines and online communities, they have advantages from a development perspective – on the one hand, LLMs do provide more accurate healthcare information; on the other hand, LLMs can help users get the health information they need in a shorter time and a concise manner. As such, using LLMs to obtain health information is a better choice compared to other information channels by many, or at least an effective complement to other methods. The advantages of LLMs stand out particularly when the public wants to obtain daily information or when they are mildly unwell – health Q&A afforded by LLMs can already help them cope with most health conditions in life. When it comes to more serious symptoms or illnesses, a thorough examination by doctors is still necessary. In addition, LLMs only provide relatively simple and error-free answers. If users want to develop professional health plans, such as making fitness plans or diet plans based on their own conditions, they still need to seek advice from professionals.
@@ 5.2. LLMs and The Doctor-Patient Relationship
Many survey participants did not agree that self-diagnosis could help doctors better judge their conditions. However, during the interviews, almost all participants admitted that after self-diagnosis, they could better communicate with doctors and enable doctors to better understand their physical conditions. We believe that everyone is the first person responsible for their own health (Resnik, 2007). If one has a good understanding of their own health status and living habits and can express them to doctors more accurately, the doctors can better judge the patient’s conditions. Therefore, LLMs’ role in helping patients conduct preliminary self-diagnosis before seeing a doctor is promising – this is a sign of being responsible for one’s own health.
@@
The issue of whether doctors should use LLMs to diagnose patients has generated significant controversy, both in surveys and in interviews. On the one hand, using LLMs for diagnosis made doctors appear unprofessional, according to the participants. On the other hand, doctors have their own limitations and cannot make accurate judgments at all times. For example, in China, medical resources are limited and doctors in less socioeconomically developed areas may not have sufficient professional knowledge and expertise (Ding et al., 2020). It is not uncommon for doctors to be physically harmed or even killed due to unsatisfactory treatment outcomes (He and Qian, 2016). Therefore, search engines or LLMs may be an important aid for doctors to make diagnoses or treatment plans. As several participants suggested, doctors can use LLMs for diagnosis, but they need to first communicate with the patients about their use and obtain their agreement. Further, doctors should apply their own judgment to LLM-generated output and explain their thought processes to the patients to build trust and ensure reliable medical outcomes.
@@
Most participants approved doctors of using LLMs to do auxiliary work such as writing medical records. Such LLM use reduces the workload of doctors and improves their work efficiency. Again, patients want to reserve the rights to know how LLM is used in the process; if LLMs generate responses that contain errors, doctors should correct them in time to avoid mistakes. As we can see, transparency and accountability (Kilhoffer et al., 2023) are important factors in LLM-assisted medical practices. Future research should consider documenting ethical incidents and errors when using LLMs for healthcare to help in this regard (Wei and Zhou, 2022).
5.3. The Future of Public Health and LLMs
With the improvement of education level and health awareness of the public, their demand for health also becomes higher. In the future, the public and doctors may not only need to diagnose diseases, but also need to have a comprehensive understanding of the cause of the disease, treatment options, mental support, and so on. As such, there is a reasonable need to combine LLMs with academic search engines such as Google Scholar to achieve such goals. For rare diseases or cutting-edge knowledge that should be held to scrutiny, LLMs can provide patients or doctors with the necessary information to make informed judgments and decisions. Prior research pointed out that ChatGPT may cite false medical literature (Frosolini et al., 2023), which remained a challenge when helping the public better retrieve health information.

### v2

@@ 3.2. Interviews
The interviews were semi-structured. We started by asking about the basic background of the participants. We divided the interview questions into three parts: (1) motivations for using LLMs for healthcare information; (2) misinformation and privacy issues encountered when using LLMs; and (3) the role of LLMs in the doctor-patient relationship. A detailed list of questions is provided in the appendix.
After each interview, we held a brief meeting to discuss emerging themes. Two authors independently conducted open and axial coding of the interview transcripts (strauss1990qualitative), and regularly met to compare their results and discuss the coding process. If the two coders could not reach a consensus, a third researcher would join the discussion to help make a final decision. In the paper, we use anonymized quotes to illustrate the main themes.
@@ 4.1. How The Public Obtain Healthcare Information? (RQ1)
LLMs are gaining popularity as a means of healthcare information seeking. Among the 219 valid surveys, 214 participants had used the Internet (including but not limited to search engines, online communities, AI-based healthcare chatbots, and LLMs) to search for healthcare information. 208 (97%) participants had searched for healthcare information from search engines, 178 (83%) participants had searched for healthcare information from online communities, and 166 (78%) participants had obtained healthcare information from LLMs. This shows that search engines are still the most mainstream way to obtain healthcare information. LLMs like ChatGPT have only been around for about a year, but they have already become a frequently used way for the public to obtain health information. In contrast, only 91 (43%) participants had used Al-led healthcare chatbots, which have existed for a longer time than LLMs. The percentage of users who used chatbots for healthcare Q&A was slightly larger than in a previous study (Nadarzynski et al., 2019), but chatbots were still not a mainstream way for the public to obtain healthcare information. We confirmed this during the interviews.
@@
There was more than one way to obtain healthcare information from the Internet, and they were often used together for better information outcomes. Of the 166 participants who had used LLMs to obtain health information, 144 participants cross-validated the healthcare information obtained from LLMs with healthcare information obtained from other sources. This suggests that LLMs have a complementary relationship with other ways of obtaining healthcare information in the digital niche. During the interviews, we found that due to the popularity of online consultation platforms in China (Ding et al., 2020; Hao, 2015; Hao et al., 2017), Some participants (P4, P6) originally from China had heavily drawn on them for healthcare information. The students majoring in medicine (P3, P16) would also use more professional medical databases, such as the Merck Manual of Diagnosis and Therapy. Other findings were largely consistent with the survey – the participants interviewed had all used search engines to search for healthcare information before, and most had obtained health information through online communities. Among the 17 participants who had used LLMs to search for health information, only P5 and P9 had used AI-led chatbots, this also shows that the popularity of AI-led Chatbots is not high.
@@
In the survey, based on some previous studies (Rice, 2006; Nadarzynski et al., 2019; Xiao et al., 2014), we roughly divided the healthcare information obtained into four categories: (1) daily life health information (such as reasonable diet); (2) medication information (such as drugs and hospitalization); (3) symptom self-diagnosis (such as searching for disease); and (4) searching for health information again after being diagnosed by a doctor due to doubts about the treatment plan. Among the 214 participants who had searched for health information, 194 (91%), 197 (92%), 190 (89%), and 176 participants had searched for the above four types of information respectively. The search frequency for the four types of information was all very high and there was no significant difference. This phenomenon showed that the public was not only paying attention to diseases but also their healthcare in daily life. Similarly, during the interviews, we found that the participants mostly used ChatGPT to search for symptom-related information, but also used it to search for daily healthcare information such as reasonable diets and fitness plans. For example, P2 only uses Chatgpt to initially develop a fitness plan:
@@
In terms of frequency of searching for health information, 121 participants in the survey searched for health information once a month or more often; during the interview, most participants said they would search for health information every one to two weeks. A few interview participants like P1 searched for healthcare information nearly every day.
@@ 4.2. Motivation for Using LLMs for Healthcare Information (RQ2)
During the interviews, we asked the participants in detail about their motivations for using LLMs to obtain health information. We were particularly curious about why participants chose to use LLMs like ChatGPT to obtain healthcare information rather than other methods such as search engines, online communities, or consulting doctors online and offline.
Compared with search engines, many interview participants used ChatGPT to obtain healthcare information because LLMs are faster and more accurate. Most of the participants were students. Before using ChatGPT to search for health information, they had used it to search for solutions to study problems. They believed that ChatGPT was more helpful to them than search engines in solving academic problems. This prompted them to start using ChatGPT for healthcare information as well.
@@
P5 also had the same idea. She believed that some subjective feelings of drug use could not be described in detail by Chatgpt. ChatGPT rarely provided emotional support. In other words, users have different orientations toward obtaining health information from LLMs and online communities.
The participants thought obtaining healthcare information from LLMs was more convenient and quick than offline or online medical consultations with doctors. Making an appointment with a doctor offline not only costs money but also takes a lot of time; even if one made an appointment with a doctor online, there was no guarantee that the doctor would always be online to answer questions. However, all the participants acknowledged that if they felt their condition was serious and needed medical equipment for examination, they would still choose to go to the hospital. In other words, all participants would only use LLMs when they were experiencing minor discomfort or searching for daily healthcare information.
From the motivations of the interview participants, we can further infer the niche of LLMs in healthcare information seeking. In general, although Chatgpt is a ”chatting” robot, most users who use Chatgpt to search for medical and health information, regard Chatgpt (or other LLMs) as a supplement or substitute for search engines. In other words, LLMs have competitive and complementary relationships with search engines and have a complementary relationship with online consultation and online communities. Users expect LLMs and search engines to provide accurate medical advice; expect online communities to provide mental support and expect online medical consultations to provide professional medical services across spatial limitations.
@@ 4.3.1. Misinformation and Inaccuracy of LLM Responses
LLMs were deemed more accurate than other information-seeking channels for healthcare information. In the survey, we designed two Likert-scale questions to probe this issue: “You can find more accurate health information with LLMs (like ChatGPT) than search engines or online communities” and “Compared with using search engines or online communities, using LLMs can reduce misinformation.” For the first statement, 109 (51%) participants “agree” or “strongly agree” with it, and 36 (17%) participants “disagree” or “strongly disagree” with it. For the second statement, 111 (52%) participants “agree” or “strongly agree” with it, and 51 (24%) participants “disagree” or “strongly disagree” with it. As we can see, more than half of the participants agreed that LLMs are more accurate and can produce less misinformation than other methods like search engines.
@@
During the interviews, we also asked participants how they judged misinformation, or how they prevented “hallucinations” from LLMs. Most of the participants did not completely give up using search engines or online communities, and they would cross-verify the health information obtained from LLMs with health information obtained through other channels. Some participants, such as P12 and P14, would consult doctors after asking LLMs for health information, and their trust in LLMs was enhanced by the doctor’s feedback. However, there are exceptions. P8 said that since using Chatgpt, he has never used a search engine again because he hates the misinformation on Google.
@@ 4.3.2. Doctor-Patient Relationship
In the survey, we designed three questions to explore this question: “Patients using LLMs to self-diagnose before asking the doctor can improve the doctor’s diagnostic performance,” “I think doctors using LLMs to diagnose and treat conditions is reasonable” and “I think doctors using LLMs for medical assistance work (such as writing medical records) is reasonable.” For the first statement, 120 (56%) participants “agree” or “strongly agree” with the statement, and 42 (20%) participants “disagree” or “strongly disagree” with the statement. For the second statement, 106 (50%) participants “agree” or “strongly agree” with the statement, and 47 (22%) participants “disagree” or “strongly disagree” with the statement. For the last statement, 140 (65%) participants “agree” or “strongly agree” with the statement, and 30 (14%) participants “disagree” or “strongly disagree” with the statement. It can be seen that most people thought it reasonable for doctors to use LLMs to perform auxiliary work such as writing medical records. When it came to using LLMs for diagnosis or self-diagnosis, the participants were more doubtful. Most participants also supported that the use of LLMs for self-diagnosis can help doctors make better diagnoses.
During the interviews, most of our participants believed that self-diagnosis assisted by LLMs could help doctors better judge their conditions. This is consistent with the survey. P6 talked about the importance of knowing one’s own medical history and physical conditions before seeing a doctor:
@@
However, a few participants believed that LLMs could be used for diagnosis since AI was an unstoppable trend and would eventually replace doctors. P7 noted:
@@ 5. Discussion
In this section, we discuss recommendations for the public to use LLMs to obtain health information and conduct self-diagnosis and recommendations for doctors to use LLMs to diagnose patients. We further give recommendations for future research in medical LLMs and give inventions for the future development direction of medical LLMs.
@@ 5.1. Recommendations for The Public
Since everyone’s physical condition is different, the same symptoms may come from different diseases for different people. We ought to fully consider these medical backgrounds and make contextual recommendations for the public regarding using LLMs to obtain health information to maximize public interests.
According to our results, although LLMs cannot completely solve the problem of misinformation generated by search engines and online communities, they have advantages from a development perspective – on the one hand, LLMs do provide more accurate healthcare information; on the other hand, LLMs can help users get the health information they need in a shorter time and a concise manner. As such, using LLMs to obtain health information is an effective complement to other information channels. The advantages of LLMs stand out particularly when the public wants to obtain daily information or when they are mildly unwell – health Q&A afforded by LLMs can already help them cope with most health conditions in life. But we also need to point out, LLMs only provide relatively simple and error-free answers. For serious and complex diseases, or when users need to develop personalized fitness/weight loss programs for themselves, LLMs are powerless. When it comes to more serious symptoms or illnesses, a thorough examination by doctors is still necessary. Of particular concern are ”hallucinations”: some patients believe that they do not have the disease, so they constantly enter ”prompts” denying the existence of the disease to Chatgpt. Chatgpt will also output content that patients want to see ”there is no disease” following the user’s ideas. Patients may therefore delay treatment (Saenger et al., 2024). Similarly, LLMs still cannot be personalized perfectly today. Due to individual differences, if users want to develop professional health plans, such as fitness plans or diet plans based on their conditions, they still need to seek advice from professionals.
@@ 5.2. LLMs and The Doctor-Patient Relationship
Many survey participants did not agree that self-diagnosis could help doctors better judge their conditions. However, during the interviews, almost all participants admitted that after self-diagnosis, they could better communicate with doctors and enable doctors to better understand their physical conditions. We believe that everyone is the first person responsible for their own health (Resnik, 2007). If one has a good understanding of their own health status and living habits and can express them to doctors more accurately, the doctors can better judge the patient’s conditions. Therefore, LLMs’ role in helping patients conduct preliminary self-diagnosis before seeing a doctor is promising – this is a sign of being responsible for one’s health.
@@
The issue of whether doctors should use LLMs to diagnose patients has generated significant controversy, both in surveys and in interviews. On the one hand, using LLMs for diagnosis made doctors appear unprofessional, according to the participants. On the other hand, doctors have their limitations and cannot make accurate judgments at all times. For example, medical resources are limited and doctors in less socioeconomically developed areas may not have sufficient professional knowledge and expertise (Ding et al., 2020; Haltaufderheide and Ranisch, 2024). It is not uncommon for doctors to be physically harmed or even killed due to unsatisfactory treatment outcomes (He and Qian, 2016). There are also some doctors spreading misinformation during the COVID-19 epidemic (Sule et al., 2023). Therefore, search engines or LLMs may be an important aid for doctors to make diagnoses or treatment plans. As several participants suggested, doctors can use LLMs for diagnosis, but they need to first communicate with the patients about their use and obtain their agreement. Further, doctors should apply their judgment to LLM-generated output and explain their thought processes to the patients to build trust and ensure reliable medical outcomes.
@@
Most participants approved doctors of using LLMs to do auxiliary work such as writing medical records. Such LLM use reduces the workload of doctors and improves their work efficiency. Again, patients want to reserve the right to know how LLM is used in the process; if LLMs generate responses that contain errors, doctors should correct them in time to avoid mistakes. As we can see, transparency and accountability (Kilhoffer et al., 2023) are important factors in LLM-assisted medical practices. Future research should consider documenting ethical incidents and errors when using LLMs for healthcare to help in this regard (Wei and Zhou, 2022).
5.3. The Future of Public Health and Trusted Healthcare LLMs
From the perspective of digital niche, most users expect LLMs to be more than just a chatbot when seeking healthcare information, but a tool that replaces search engines. Users expect to obtain healthcare information accurately and timely through LLMs, which means that people have similar expectations of LLMs as professional doctors. If LLMs are expected to do the jobs of a doctor, the first thing we need to discuss is what scientific and ethical principles a trustworthy doctor should abide by. Among these, the public is particularly concerned about medical misinformation. This has triggered a scientific and ethical debate: Is the information provided by doctors ”misinformation”? How can we identify ”misinformation” in medical treatment? Medicine and Healthcare are rapidly developing – academic results that are correct today may be overturned or proven to have certain limitations tomorrow (Singhal et al., 2023a). The medical and healthcare field is a rapidly developing field, and many cutting-edge studies may contain errors or even academic fraud, such as cardiac stem cells (Knoepfler, 2018). So we must have a clear definition of misinformation in medicine and healthcare, otherwise there will be confusion. In June 2022, in order to curb the spread of misinformation, the American Medical Association adopted a policy outlining specific steps to identify and curb false and misleading medical information (Ethics and Committee, 2022). Based on this guidance, we provide some inspiration for future trustworthy medical LLMs.
Timely and rapid updating of the database (corpus) is critical. In the guidance, medical misinformation is defined as: ”Health-related information or claims that are false, inaccurate or misleading, according to the best available scientific evidence at the time”. We cannot ask doctors to answer medical questions that cannot be explained by current evidence, nor can we ask doctors to diagnose major diseases without any mistakes. The same is true for LLMs. However, our scientific research needs that LLMs can always follow the latest scientific research progress. Today, with the rapid progress of scientific research, only by constantly updating the database can LLMs ensure that they can provide users with the most accurate information based on current scientific evidence.
Effective methods of identifying misinformation are needed. For the field of medicine and health, in addition to misinformation on social media or generated by other LLMs, a more unique challenge is identifying misinformation in the literature. If LLMs are required to provide information that is most consistent with current scientific evidence, it is inevitable to incorporate the latest research literature into the database. However, a survey showed that of the more than 5,000 neuroscience articles published in 2020, 34% may be plagiarized or made up. In medicine, this proportion is 24% (Brainard, 2023). Prior research pointed out that ChatGPT may cite false medical literature (Frosolini et al., 2023) In an environment where LLMs have shifted from model-driven to data-driven (He et al., 2023), the quality of data determines the accuracy of the model. Filtering out harmful articles is of great significance to improve the accuracy of LLMs.
Balancing autonomy with medical advice. Doctors indeed need to abide by the principle of autonomy and respect patients’ right to choose their medical solutions, but this does not mean that doctors should excessively cater to patients’ ideas and make patients choose wrong or outdated diagnoses and treatment options. A good doctor should also be able to persuade patients to accept the correct diagnosis and treatment plan. This puts forward higher requirements for the dialogue system of LLMs. While avoiding hallucinations, LLMs need to use humane methods and language to help and even persuade patients to choose the best treatment options.
In the field of daily health, such as weight loss and fitness, personalized health care LLMs are a direction worthy of attention. Current LLMs still fail to develop personalized solutions based on user needs  (Yang et al., 2024). The ultimate goal of disease diagnosis and treatment is to restore patients to health, but users who use LLMs for daily health activities are likely to have different needs. Thus, understanding user needs and designing personalized LLMs become increasingly important.
Combining LLMs with search engines. Although our research found that LLMs and search engines compete, this does not mean that the two do not have a complementary relationship. A study has shown that users prefer to use search engines to search for direct, fact-based queries, and use LLMs to search for complex natural language processing problems (Caramancion, 2024). Many works combine LLMs with search engines and point out that LLMs can be excellent search engines (Ziems et al., 2023; Vu et al., 2023). This shows that integrating the advantages of both to facilitate users to search more conveniently will be a future trend.